OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 6 square root of 2 _____a0cm.?

OpenStudy (anonymous):

\[(6\sqrt{2})^2=x^2+x^2=2x^2\]